BRAINIAC:

Improper Integral Type 1

BRAINIAC:

http://prntscr.com/km2do9

mhchen:

Do you know how to integrate \[\int\limits_{1}^{a}\frac{1}{x}dx\]

mhchen:

You basically just integrate it normally, then once you have your answer, you do \[\lim_{a \rightarrow \infty} \] to see if the limit is a number. If it's infinity, then it diverges.

BRAINIAC:

yeap ^ Thank you @mhchen ;)